# Beyond Ensemble Averages: Leveraging Climate Model Ensembles for Subseasonal Forecasting

### v3

@@ 8.1 The efficacy of machine learning for SSF
Several hypotheses might explain why ML is an effective approach for SSF, and we probe those hypotheses in this section. Appendix C.4 provides an illustration of precipitation predictions using different methods, demonstrating how these predictions vary spatially.
@@
Past works use ensemble mean as an input feature to machine learning methods in addition to the climate variables [12, 37]. Ensembles provide valuable information not only about expected climate behavior, but also variance or uncertainty in multiple dimensions; methods that rely solely on ensemble mean lack information about this variance. The dynamical models that underpin each ensemble member may have systematic errors, either in the mean, the variability, or conditioned upon certain elements of the initial conditions that are not readily apparent to the forecast model. The more recently initialized an ensemble member is, the better it usually performs. While taking the average of these ensemble members may cancel out the deficiencies of each individual member, it is also possible that details of the systematic errors of each may be directly discovered and corrected by a machine learning model independently. Therefore, using a single ensemble statistic, such as the ensemble mean, as a feature may not fully capitalize on the information provided by using all the ensemble forecasts as features.
@@
In our experiments, we find that using all available ensemble members enhances the prediction quality of our approaches. As an illustration, we show the results of the LR, RF, U-Net, and the stacked model trained on all ensemble members, compared to the ML models trained on the ensemble mean only. In addition to the full ensemble or the ensemble mean, we use other available features (as in our previous regression results). Table 7 and Table 8 demonstrate the precipitation and temperature forecasting results. For the linear regression, utilizing the ensemble mean with all other features produces the best test performance compared to using the full ensemble. Such behavior is not surprising for the LR since the full ensemble incorporates large variance across ensemble members, which may result in a worse linear fit.For the U-Net, RF, and stacked model, we observe significant performance improvement in terms of having at least one standard error smaller MSE, when using the full ensemble instead of using the ensemble mean. When we compare the performance of learned models using only the ensemble mean to that of learned models that use both the ensemble mean and the ensemble standard deviation for each spatial location, we find that the addition of the standard deviation feature does not provide enough information to significantly improve the performance of ML models, and in fact the U-Net that exhibits a performance degradation – a potential a sign of overfitting. These observations are visually supported by Figure 5 and Figure 6, where the $R^{2}$ heatmaps of our methods (except the U-Net) utilizing ensemble mean and standard deviation closely resemble the performance of methods solely relying on ensemble mean. We conclude that the full ensemble contains important information for SSF aside from the ensemble mean, and our models can capitalize on this information for precipitation and temperature forecasting.
@@
Table 7: Precipitation forecasting performance comparison of the LR, RF, U-Net, and stacked model trained using the ensemble mean, using the sorted ensemble members or using the original ensemble, in addition to other features. Scores on the test data are reported and NCEP-CFSv2 data is used. The best results are in bold. MSE is reported in mm.
@@
Figure 5: Precipitation regression test $R^{2}$ heatmaps of LR, U-Net, RF and stacked model trained using ensemble mean only, using sorted and shuffled ensemble, or using the full ensemble. The NCEP-CFSv2 ensemble is used. See Section 8.1 for details.
Table 8: Temperature forecasting performance comparison of the LR, RF, U-Net, and stacked model trained using the ensemble mean, using the sorted ensemble members or using the original ensemble, in addition to other features. Scores on the test data are reported and NCEP-CFSv2 data is used. The best results are in bold. MSE is reported in ${}^{\circ}C$.
@@
Shuffled ensemble + all features -0.03 3.95$\pm$0.03
@@
U-Net Ensemble mean & std + all features -0.19 4.61$\pm$0.03
Shuffled ensemble + all features -0.29 4.75$\pm$0.03
@@
We can perform a statistical test to verify that the performance discrepancies between using the ensemble mean and using the full ensemble are statistically significant for the stacked model. As before, let $\hat{y}_{t,l}$ refer to the estimate under our usual stacked model (i.e., with all ensemble members). Let $\hat{y}_{t,l}^{\text{SEA}}$ refer to a stacked model with just the ensemble mean in place of each ensemble member. We can employ a sign test framework [61, 62] to compare model performance under minimal distributional assumptions. Namely, we only make the following i.i.d. assumption over the time dimension:
[…]
$H_{0,l}:p_{l}=0.5,\;\;H_{0,l}:p_{l}>0.5$
Thus, our overall test for significance is for the global null hypothesis $H_{0}=\cap_{l=1}^{3274}H_{0,l}$. Specifically, we calculate a p-value for each $H_{0,l}$, and then we check whether any of these p-values is below a Bonferroni-corrected threshold of $0.05/3274=1.53\times 10^{-5}$, where 3274 refers to the number of locations. In fact, the minimum p-values for this test with precipitation and regression alike are far below this threshold ($1.68\times 10^{-10}$ and $4.42\times 10^{-24}$, respectively). This allows us to reject the global null hypothesis for both temperature and precipitation, and we conclude that including the full ensemble in our stacked model significantly outperforms including just the ensemble mean.
@@
We consider the hypothesis that there is a set of $k$ ensemble members that are always best. To test this hypothesis, we use a training period to identify which $k$ members perform best for each location, and then during the test period, compute the average of only these $k$ ensemble members. The performance of this approach depends on $k$, the number of ensemble members we allow to be designated “good,” but the performance for any $k$ never exhibited a significant improvement over the ensemble mean.
@@
If the ensemble members have different levels of accuracy over various seasons, locations, and conditions, then a machine learning model may be learning when to “trust” each member. We know that our ensemble members are lagged, meaning they are initialized at different times. We believe that each ensemble member encapsulates valuable information derived from the underlying physical model during each initialization. To investigate the impact of ensemble member ordering, we perform the following experiment: we randomly permute ensemble members at every time step $t$ for all locations (preserving the spatial information), and apply our ML models to these shuffled ensembles. From Table 7 and Table 8, this approach negatively affects the performance of the ML models compared to using the full ensemble with the original order. Random shuffling disrupts the inherent temporal patterns specific to each ensemble member. One possible explanation is that the learned models lose the ability to capture temporal dependencies associated with the original ensemble order, even if spatial information remains intact.
@@
In addition, we train and evaluate our ML models with the sorted forecasts as the features and compare them to the results, which used the original ordering of the ensemble members. The ensemble members are sorted independently for each latitute/longitude pair, disrupting the spatial structure. According to Table 7 and Table 8, using the sorted ensemble drastically degrades U-Net’s performance, which is essentially because we sort the ensemble members for each location individually, and sorting the ensemble members individually for each location may break the spatial correlations that the U-Net is learning. In the case of precipitation regression with the stacking model from Table 7, the MSE of the sorted approach is 2.24, which is worse than the 2.07 MSE for using the original ordering. In the case of temperature forecasting from Table 8, the MSE of the sorted approach is 3.70, which is much worse than the 3.11 MSE for using the original ordering. The mean $R^{2}$ of the sorted approach is also lower compared to the original ordering. In both cases, the performance is better when we feed in the features in such a way that the machine learning model has an opportunity to learn aspects of each ensemble member, not merely their order statistics. Therefore, imposing a symmetric treatment of ensemble members degrades performance. Figure 5 and Figure 6 shows the corresponding $R^{2}$ heatmaps of our models for precipitation and temperature regression tasks.
@@ 8.2 Using spatial data
There are a number of ways to incorporate information about location in our models. U-Net has access to the spatial dependencies by its design. Specifically, our U-Net inputs the spatial location of each point in the map. Naively, we might represent each location using the latitude and longitude values. Alternatively, we may use positional encoding, which is known to be beneficial in many ML areas, not only in NLP (as we mention in Section 6.1). PE captures the order (or position) and allows one to learn the contextual relationships (local context – relationships between nearby elements, and global context dependencies across the entire sequence). We assume that the PE approach represents spatial information in a manner more accessible to our learned models.
As an illustration, Table 9 and Figure 7 demonstrates the performance of a stacked model using LR, RF and U-Net trained using positional encoding, using latitude and longitude values and using no features representing the spatial information (no PE and no latitude or longitude values). Other inputs to the LR, RF and U-Net models are ensemble member forecasts, lagged target variable, climate variables, and SSTs (no SSTs in the case of temperature forecasting). The results suggest that using PE enhances the predictive skill of our models, compared to using just the lat/lon values or no location information, especially for the temperature forecasting task. Using no information about locations hurts the performance for precipitation regression. Thus, our models can account for spatial dependencies using input features, and PE is more beneficial than the raw latitude and longitude information. These findings on PE effectiveness are consistent with prior findings in ML. For example, [63] investigate the efficacy of PE in the context of the visual transformer model used for image classification and object detection. We show a more detailed analysis with results for the LR, RF and U-Net in Appendix C.2.
Table 9: Test performance comparison of the stacked model of LR, RF, and U-Net trained using no spatial features, using latitude and longitude values or using PE. Utilizing spatial representations, including PE, latitude, and longitude values, helps advance the predictive skill. Furthermore, using positional encoding is more beneficial than using raw latitude and longitude values. The best results are in bold. MSE is reported in mm for precipitation and in ${}^{\circ}C$ for temperature.
@@ 8.3 Variable importance
One consideration when implementing ML for SSF is that ML models can incorporate side information (such as spatial information, lagged temperature and precipitation values, and climate variables). We explore the importance of the various components of side information in this section. We see that including the observational climate variables improve the performance for both the random forest and the U-Net when doing precipitation regression. Furthermore, including positional encoding of the locations improves the performance of the U-Net, while the principle components of the sea surface temperature do not make notable difference in the case of temperature prediction.
Table 10: Grouped feature importance results on validation for precipitation regression task using NCEP-CFSv2 ensemble members. The results suggest that using additional observational information helps to improve the performance of learning-based models for this task. –"– means a repetition of features that are used above. For example, in the U-Net part of the table, “–"– & lags” means that ensemble members, PE, and lags are used as features and “–"– & SSTs” means ensemble members, PE and lags, land features and SSTs are used as features. The best results are in bold. MSE is reported in mm.
@@
More specifically, Table 10 summarizes grouped feature importance of precipitation regression using the NCEP-CFSv2 ensemble. We observe that models, in particular random forest and U-Net, trained on all available data achieve the best performance. In the case of linear regression, the SSTs features are neither very helpful nor actively harmful. Therefore, in order to be consistent, we use predictions of these models trained on all features as input to the stacking model.
Table 11: Grouped feature importance results on validation for temperature regression task using NCEP-CFSv2 ensemble members. The results demonstrate that using some additional information may yield only very small improvements in predictive skill, and in some cases the side information may decrease predictive skill. –"– means a repetition of features that are used above. The best results are in bold. MSE is reported in ${}^{\circ}C$.
@@
Table 11 summarizes grouped feature importance of temperature regression using the NCEP-CFSv2 ensemble. In this case, adding some types of side information may yield only very small improvements to predictive skill, and in some cases the additional information may decrease predictive skill. On the one hand, this effect can be explained by different training set sizes for different models: as we outline in Section 5.4, the training set size for RF is $n=TL$, while $n=T$ for U-Net. This effect also may be a sign of overfitting, as temperature forecasting presents a comparatively less complex challenge than precipitation forecasting. We also note that SSTs provide only marginal (if any) improvement in predictive skill, in part because Pacific SSTs are less helpful away from the western U.S. [64, 65]. It could also be that information from the SSTs is already being well-captured by the output from the dynamical models, and thus including observed SSTs is not providing much additional information. In order to be consistent, we use predictions of these models trained on all features except SSTs as an input to the stacking model.
@@ 9 Conclusions and future directions
Our results also suggest several important directions of future research. In terms of features, there are many climate forecasting ensembles computed by organizations such as NOAA and ECMWF. This paper focuses on ensembles in which ensemble members have a distinct ordering (in terms of lagged initial conditions used to generate them), but other ensembles correspond to initial conditions or parameters drawn independently from some distribution. Leveraging such ensemble forecasts, and potentially jointly leveraging ensemble members from multiple distinct ensembles, may further improve the predictive accuracy of our methods.
In terms of models, new neural architecture models such as transformers have shown remarkable performance on a number of image analysis tasks [66, 67, 68, 59] and have potential in the context of forecasting climate temperature and precipitation maps. Careful study is needed, as past image analysis work using transformers generally uses large quantities of training data, exceeding what is available in SSF contexts.
In terms of methods, two outstanding challenges are particularly salient to the SSF community. The first is uncertainty quantification; that is, we wish not only to forecast temperature or precipitation, but also to predict the likelihood of certain extreme events. Our work on quantile regression is an important step in this direction, and statistical methods like conformal quantile regression [69] may provide additional insights. Second, we see in Figure 23 that, at least in some geographic regions, the distribution of ensemble hindcast and forecast data may be quite different. Employing methods that are more robust to distribution drift [70, 71, 72] is particularly important not only to handling forecast and hindcast data, but also for accurate SSF in a changing climate.
Acknowledgements

### v4

@@ 8.1 The efficacy of machine learning for SSF
Several hypotheses might explain why ML may be a promising approach for SSF, and we probe those hypotheses in this section.
@@
Past works use ensemble mean as an input feature to machine learning methods in addition to the climate variables [12, 37]. Ensembles provide valuable information not only about expected climate behavior but also variance or uncertainty in multiple dimensions; methods that rely solely on ensemble mean lack information about this variance. Ensemble members may have systematic errors, either in the mean or the variability, arising from different initial conditions of the corresponding dynamic model that are not readily apparent to users. The more recently initialized an ensemble member is, the better it usually performs. While taking the average of these ensemble members may cancel out the deficiencies of each individual member, it is also possible that details of each member’s systematic errors may be directly discovered and corrected independently by a machine learning model. Therefore, using a single ensemble statistic, such as the ensemble mean, as a feature may not fully capitalize on the information provided by using all members of the lagged ensemble as features.
@@
In our experiments, we find that using all available ensemble members enhances the prediction quality of our approaches. As an illustration, we show the results of the LR, RF, U-Net, and the stacked model trained on all ensemble members, compared to the ML models trained on the ensemble mean only. In addition to the full ensemble or the ensemble mean, we use other available features (as in our previous regression results). Table 7 and Table 8 demonstrate the precipitation and temperature forecasting results. For the linear regression, utilizing the ensemble mean with all other features produces the best test performance compared to using the full ensemble. Such behavior is not surprising for the LR since the full ensemble incorporates large variance across ensemble members, which may result in a worse linear fit. For the U-Net, RF, and stacked model, we observe significant performance improvements, in terms of having at least one standard error smaller MSE, when using the full ensemble instead of using the ensemble mean. When we compare the performance of learned models using only the ensemble mean to that of the learned models that use both the ensemble mean and the ensemble standard deviation for each spatial location, we find that the addition of the standard deviation feature does not provide enough information to significantly improve the performance of ML models, and in fact the U-Net that exhibits a performance degradation – a potential a sign of overfitting. These observations are visually supported by Figure 5 and Figure 6, where the $R^{2}$ heatmaps of our methods (except the U-Net) utilizing ensemble mean and standard deviation closely resemble the performance of methods solely relying on ensemble mean. We conclude that the full ensemble contains important information for SSF aside from the ensemble mean, and our models can capitalize on this information for precipitation and temperature forecasting.
We can perform a statistical test to verify that the performance discrepancies between using the ensemble mean and using the full ensemble are statistically significant for the stacked model. As before, let $\hat{y}_{t,l}$ refer to the estimate under our usual stacked model (i.e., with all ensemble members). Let $\hat{y}_{t,l}^{\text{SEA}}$ refer to a stacked model with just the ensemble mean as a feature, instead of all ensemble members. We can employ a sign test framework [61, 62] to compare model performance under minimal distributional assumptions. Namely, we only make the following i.i.d. assumption over the time dimension:
[…]
$H_{0,l}:p_{l}=0.5,\;\;H_{0,l}:p_{l}>0.5$
Thus, our overall test for significance is for the global null hypothesis $H_{0}=\cap_{l=1}^{3274}H_{0,l}$. We calculate a p-value for each $H_{0,l}$, and then we check whether any of these p-values is below a Bonferroni-corrected threshold of $0.05/3274=1.53\times 10^{-5}$, where 3274 refers to the number of locations. In fact, the minimum p-values for this test with precipitation and regression alike are far below this threshold ($1.68\times 10^{-10}$ and $4.42\times 10^{-24}$, respectively). This allows us to reject the global null hypothesis for both temperature and precipitation, and we conclude that including the full ensemble in our stacked model significantly outperforms including just the ensemble mean.
@@
Table 7: Precipitation forecasting performance comparison of the LR, RF, U-Net, and stacked model trained using the ensemble mean, using the sorted ensemble members, or using the original ensemble, in addition to other features. Scores on the test data are reported, and NCEP-CFSv2 data is used. The best results are in bold. MSE is reported in squared mm.
@@
Figure 5: Precipitation regression test $R^{2}$ heatmaps of LR, U-Net, RF, and stacked model trained using ensemble mean only, using sorted and shuffled ensemble, or using the full ensemble. The NCEP-CFSv2 ensemble is used. See Section 8.1 for details.
Table 8: Temperature forecasting performance comparison of the LR, RF, U-Net, and stacked model trained using the ensemble mean, using the sorted ensemble members, or using the original ensemble, in addition to other features. Scores on the test data are reported and NCEP-CFSv2 data is used. The best results are in bold. MSE is reported in squared ${}^{\circ}C$.
@@
Shuffled ensemble + all features 0.03 3.95$\pm$0.03
@@
U-Net Ensemble mean & std + all features 0.19 4.61$\pm$0.03
Shuffled ensemble + all features 0.29 4.75$\pm$0.03
@@
We consider the hypothesis that there is a set of $k$ ensemble members that are always best. To test this hypothesis, we use a training period to identify which $k$ members perform best for each location, and then during the test period, compute the average of only these $k$ ensemble members. The performance of this approach depends on $k$, the number of ensemble members we allow to be designated “good.” We have not found that the performance for any $k$ exhibits a significant improvement over the ensemble mean.
@@
If the ensemble members have different levels of accuracy over various seasons, locations, and conditions, then a machine learning model may be learning when to “trust” each member. We know that our ensemble members are lagged, meaning they are initialized at different times. We believe each ensemble member encapsulates valuable information derived from the underlying physical model during each initialization. To investigate the impact of ensemble member order, we perform the following experiment: we randomly permute ensemble members at every time step $t$ for all locations (preserving the spatial information) and apply our ML models to these shuffled ensembles. From Table 7 and Table 8, this approach negatively affects the performance of the ML models compared to using the full ensemble with the original order. One possible explanation is that the learned models lose the ability to learn which ensemble member to trust, as this information is tied to the initialization time of each ensemble member. Even though the spatial information remains intact after the shuffling, the models can no longer exploit dependencies associated with the original ensemble structure.
@@
Additionally, we conduct an experiment designed to test whether it is important to keep track of which ensemble member made each prediction or whether it is the distribution of predictions that is important. The modeling approach for the former would be to feed in ensemble member 1’s forecast as the first feature, ensemble member 2’s forecast as the second feature, etc. The modeling approach under the distributional hypothesis is to make the smallest prediction be the first feature, the second-smallest prediction be the second feature, and so on – i.e., we sort the ensemble forecasts for each location separately. Note that this entails treating the ensemble members symmetrically: the model would give the same prediction if ensemble member 1 predicted $a$ and ensemble member 2 predicted $b$ or if ensemble member 1 predicted $b$ and ensemble member 2 predicted $a$. In statistical parlance, this is passing in the order statistics of the forecasts as the features rather than their original ordering. Note that for NCEP-CFSv2, ensemble forecasts are originally ordered according to the time their initial conditions are set [15]. According to Table 7 and Table 8, using the sorted ensemble drastically degrades U-Net’s performance, which is essentially because we sort the ensemble members for each location individually, and sorting the ensemble members individually for each location may hamper the ability for the U-Net to learn spatial structure. In the case of precipitation regression with the stacking model from Table 7, the MSE of the sorted approach is 2.24, which is worse than the 2.07 MSE for using the original ordering. In the case of temperature forecasting from Table 8, the MSE of the sorted approach is 3.70, which is much worse than the 3.11 MSE for using the original ordering. The mean $R^{2}$ of the sorted approach is also lower compared to the original ordering. In both cases, the performance is better when we feed in the features in such a way that the machine learning model has an opportunity to learn aspects of each ensemble member, not merely their order statistics. Therefore, imposing a symmetric treatment of ensemble members degrades performance. Figure 5 and Figure 6 shows the corresponding $R^{2}$ heatmaps of our models for precipitation and temperature regression tasks.
@@ 8.2 Using spatial data
There are several ways to incorporate information about location in our models. U-Net has access to spatial dependencies through its design. Specifically, our U-Net inputs the spatial location of each point in the map. Naively, we might represent each location using the latitude and longitude values. Alternatively, we may use positional encoding, which is known to be beneficial in many ML areas, not only in NLP (as we mention in Section 6.1). PE captures the order (or position) and allows one to learn the contextual relationships (local context – relationships between nearby elements and global context dependencies across the entire sequence). We assume that the PE approach represents spatial information in a manner more accessible to our learned models.
As an illustration, Table 9 and Figure 7 demonstrate the performance of a stacked model using LR, RF, and U-Net trained using positional encoding, using latitude and longitude values and using no features representing the spatial information (no PE and no latitude or longitude values). Other inputs to the LR, RF, and U-Net models are ensemble member forecasts, lagged target variable, climate variables, and SSTs (except in the case of temperature forecasting). The results suggest that using PE enhances the predictive skill of our models, compared to using just the lat/lon values or no location information, especially for the temperature forecasting task. Using no information about locations hurts the performance of precipitation regression. Thus, our models can account for spatial dependencies using input features, and PE is more beneficial than the raw latitude and longitude information. These findings on PE effectiveness are consistent with prior findings in ML. For example, [63] investigate the efficacy of PE in the context of the visual transformer model used for image classification and object detection. We show a more detailed analysis with results for the LR, RF, and U-Net in Appendix C.1.
Table 9: Test performance comparison of the stacked model of LR, RF, and U-Net trained using no spatial features, using latitude and longitude values, or using PE. Utilizing spatial representations, including PE, latitude, and longitude values, helps advance the predictive skill. Furthermore, using positional encoding is more beneficial than using raw latitude and longitude values. The best results are in bold. MSE is reported in squared mm for precipitation and in squared ${}^{\circ}C$ for temperature.
@@ 8.3 Variable importance
One consideration when implementing ML for SSF is that ML models can incorporate side information (such as spatial information, lagged temperature and precipitation values, and climate variables). In this section, we explore the importance of the various components of side information. We see that including the observational climate variables improves the performance of the random forest and the U-Net when doing precipitation regression. Furthermore, including positional encoding of the locations improves the performance of the U-Net, while the principle components of the sea surface temperature do not make a notable difference in the case of temperature prediction.
More specifically, Table 10 summarizes grouped feature importance of precipitation regression using the NCEP-CFSv2 ensemble. We observe that models, in particular random forest and U-Net, trained on all available data achieve the best performance. In the case of linear regression, the SSTs features are neither very helpful nor actively harmful. Therefore, to be consistent, we use predictions of these models trained on all features as input to the stacking model.
Table 10: Grouped feature importance results on validation for precipitation regression task using NCEP-CFSv2 ensemble members. The results suggest that using additional observational information helps to improve the performance of learning-based models for this task. –"– means a repetition of features that are used above. For example, in the U-Net part of the table, “–"– & lags” means that ensemble members, PE, and lags are used as features and “–"– & SSTs” means ensemble members, PE and lags, land features and SSTs are used as features. The best results are in bold. MSE is reported in squared mm.
@@
Table 11 summarizes grouped feature importance of temperature regression using the NCEP-CFSv2 ensemble. In this case, adding some types of side information may yield only very small improvements to predictive skill, and in some cases, the additional information may decrease predictive skill. On the one hand, this effect can be explained by different training set sizes for different models: as we outline in Section 5.4, the training set size for RF is $n=TL$, while $n=T$ for U-Net. This effect also may be a sign of overfitting, as temperature forecasting presents a comparatively less complex challenge than precipitation forecasting. We also note that SSTs provide only marginal (if any) improvement in predictive skill, in part because Pacific SSTs are less helpful away from the western U.S. [64, 65]. It could also be that information from the SSTs is already being well-captured by the output from the dynamical models, and thus, including observed SSTs does not provide much additional information. In order to be consistent, we use predictions of these models trained on all features except SSTs as input to the stacking model.
Table 11: Grouped feature importance results on validation for temperature regression task using NCEP-CFSv2 ensemble members. The results demonstrate that using some additional information may yield only very small improvements in predictive skill, and in some cases, the side information may decrease predictive skill. –"– means a repetition of features that are used above. For example, in the U-Net part of the table, “–"– & lags” means that ensemble members, PE, and lags are used as features and “–"– & SSTs” means ensemble members, PE and lags, land features and SSTs are used as features. The best results are in bold. MSE is reported in squared ${}^{\circ}C$.
@@ 9 Conclusions and future directions
Our results also suggest several important directions for future research. In terms of features, there are many climate forecasting ensembles computed by organizations such as NOAA and ECMWF. This paper focuses on ensembles in which ensemble members have a distinct ordering (in terms of lagged initial conditions used to generate them), but other ensembles correspond to initial conditions or parameters drawn independently from some distribution. Leveraging such ensemble forecasts and potentially jointly leveraging ensemble members from multiple distinct ensembles may further improve the predictive accuracy of our methods.
In terms of models, new neural architecture models such as transformers show remarkable performance on several image analysis tasks [66, 67, 68, 59] and have potential in the context of forecasting climate temperature and precipitation maps. A careful study is needed, as past image analysis work using transformers generally uses large quantities of training data, exceeding what is available in SSF contexts. Recent advancements in data-driven global weather forecasting models, such as Pangu-Weather [69], FourCastNet [70], and GenCast [71], demonstrate the potential of ML techniques to enhance forecasting capabilities across various timescales. These models outperform traditional numerical weather prediction approaches, suggesting that similar data-driven methods may hold promise for improving SSF quality.
In terms of methods, two outstanding challenges are particularly salient to the SSF community. The first is uncertainty quantification; that is, we wish not only to forecast temperature or precipitation but also to predict the likelihood of certain extreme events. Our work on quantile regression is an important step in this direction and statistical methods like conformalized quantile regression [72] may provide additional insights. Second, we see in Figure 23 that, at least in some geographic regions, the distribution of ensemble hindcast and forecast data may be quite different. Employing methods that are more robust to distribution drift [73, 74, 75] is particularly important not only for handling forecast and hindcast data but also for accurate SSF in a changing climate.
Acknowledgments